Circular No. 9106

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2010 A3 (HILL) R. E. Hill reports his discovery of another comet with the Catalina

0.68-m Schmidt telescope (discovery observation tabulated below); four coadded 30-s CCD exposures show a well-condensed nuclear condensation 4′′–5′′ across with a possible 2′′ coma beyond that and a broad, diffuse tail 10′′–15′′ long in p.a. 80^{o}. Four follow-up co-added 30-s CCD exposures by A. Boattini with the Mt. Lemmon 1.5-m reflector on Jan. 8.16–8.17 show a well-condensed comet with a coma ∼ 5′′–6′′ wide and a broad, fan-shaped tail ∼ 15′′ long in p.a. 75^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometrists have also noted the object’s cometary appearance. A. Novichonok, Kondopoga, Russia, reports that five stacked 300-s images taken remotely by T. Kryachko with the 0.3-m f/8 Astrotel reflector at Zelenchukskaya Station on Jan. 8.7 UT show a round, diffuse coma of diameter 10′′ and a wide, faint 15′′ tail in p.a. 280^{o}. H. Sato (Ota-ku, Tokyo, Japan, remotely using an RAS Observatory 0.25-m f/3.4 reflector near Mayhill, NM, U.S.A.; Jan. 9.1 UT) finds a 10′′ coma elongated toward p.a. 95^{o}. W. H. Ryan (Magdalena Ridge Observatory, 2.4-m f/8.9 reflector; Jan. 9.12–9.15) notes a coma and a faint tail in p.a. ∼ 80^{o} on R-band images. E. Guido (Castellammare di Stabia, Italy) reports that he and G. Sostero co-added eighteen 120-s unfiltered exposures taken remotely as done by Sato (above) on Jan. 9.14–9.16, showing a coma nearly 10′′ in diameter, slightly elongated toward p.a. 50^{o}. K. Kadota (Ageo, Saitamaken, Japan, 0.25-m f/5 reflector; Jan. 9.5) reports the comet to be diffuse with central condensation, with a coma diameter of 0′.4 and a faint tail 1′.5 long in p.a. 80^{o}.

2010 UT α_{2000} δ_{2000} Mag. Observer
Jan. 8.09596 1^{h}44^{m}01.^{s}92 +8^{o}13′15′′.6 17.3 Hill

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2010-A48.

T = 2010 Mar. 27.292 TT  ω = 38.^{o}455
Ω = 63.847 }2000.0
q = 1.65798 AU  i = 15.064

COMET 217P/2009 F3 (LINEAR) Visual total-magnitude estimates by M. Goiato (Aracatuba, Brazil,

22-cm reflector): Nov. 13.12 UT, 10.6; Dec. 6.06, 10.9:; 18.21, 12.1.

2010 January 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green